# Asymptotic-preserving semi-implicit finite volume scheme for Extended Magnetohydrodynamics

### v1

@@ Acknowledgement
This work was funded by Los Alamos National Laboratory under the grant “Algorithm/hardware/software Co-design for High Energy Density Applications,” at the University of Michigan. The authors are grateful to Jonah Miller and Patrick Mullen for valuable advise on this work.
@@ Appendix C Temperature relaxation
To test whether the electron-ion energy exchange source term is working in terms of equilibrating electron and ion temperatures, Sod shock tube (e.g., plasma $\beta\rightarrow\infty$ thus $B_{0}=0$) for Hydogen plasma is considered as follows,
[…]
where $L_{0}=10^{4}$ and $t_{0}=10^{2}$ are chosen such that the electron-ion energy exchange source term becomes significant. Two densities, $n_{0}=10^{20}\,\text{m}^{-3}$ and $n_{0}=10^{30}\,\text{m}^{-3}$, are used to test weakly and strongly collisional regimes, respectively, with the latter expected to yield $T_{e}=T_{i}$. The results in Fig. 14 show that only the dense plasma reaches ion–electron temperature equilibrium by the end of the simulation.
(a) For $n_{0}=10^{20}\text{m}^{-3}$, the post-shock electron temperature is expected to be lower than the ion temperature, since electrons experience primarily adiabatic compression heating, whereas electron–ion energy exchange remains weak in a weakly collisional plasma.
(b) At $n_{0}=10^{30}\text{m}^{-3}$ where both ions and electrons temperatures are expected to equilibrate at end time.
Figure 14: Sod shock tube problem is used to test for the electron-ion energy exchange source term at different densities.

### v2

@@ Acknowledgement
The authors are grateful to Jonah Miller, Patrick Mullen, and Luke Roberts for valuable advise on this work. This work has been assigned document release number LA-UR-25-31337. Support was provided by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
@@ Appendix C Temperature relaxation
To test whether the electron-ion energy exchange source term is working in terms of equilibrating electron and ion temperatures, Brio-Wu shock tube problem for Hydogen plasma is considered. $L_{0}=10^{7}$ and $t_{0}=10^{5}$ are chosen such that the electron-ion energy exchange source term becomes significant. Two densities, $n_{0}=10^{26}\,\text{m}^{-3}$ and $n_{0}=10^{30}\,\text{m}^{-3}$, are used to test weakly and strongly collisional regimes, respectively, with the latter expected to yield $T_{e}=T_{i}$. $B_{0}$ is chosen as 4.568 such that $\frac{B_{0}^{2}}{\mu_{0}n_{0}m_{i}v^{2}}$ go to 1. The results in Fig. 14 show that only the dense plasma reaches ion–electron temperature equilibrium by the end of the simulation.
Figure 14: At $n_{0}=10^{30}$ where both ions and electrons temperatures are expected to equilibrate to equilibrium temperature.
Figure 15: For $n_{0}=10^{26}$, the post-shock electron temperature is expected to be lower than the ion temperature, since electrons experience primarily adiabatic compression heating while electron–ion energy exchange remains weak in a weakly collisional plasma.